### v1

# Simulating non-Brownian suspensions with non-homogeneous Navier slip boundary conditions

Abstract
Fluid-structure interactions are commonly modeled using no-slip boundary conditions. However, small deviations from these conditions can significantly alter the dynamics of suspensions and particles, especially at the micro and nano scales. This work presents a robust implicit numerical method for simulating non-colloidal suspensions with non-homogeneous Navier slip boundary conditions. Our approach is based on a regularized boundary integral formulation, enabling accurate and efficient computation of hydrodynamic interactions. This makes the method well-suited for large-scale simulations. We validate the method by comparing computed drag forces on homogeneous and Janus particles with analytical results. Additionally, we consider the effective viscosity of suspensions with varying slip lengths, benchmarking against available analytical no-slip and partial-slip theories.
keywords:
Navier slip , partial slip , Stokes flows
1 Introduction
The study of suspensions dynamics often assume the no-slip boundary conditions (BCs) on the particles surfaces for practical purposes. The no-slip condition states that the fluid velocity at a solid boundary matches the velocity of the boundary itself. This approximation has been successfully used to explain many fluid dynamics phenomena, such as the dynamics of particles in flow and the rheology of suspensions [1, 2]. However, the no-slip BCs are not always satisfied, and even small deviations can significantly influence the hydrodynamics of small particles [3, 4]. The no-slip BCs can be indeed generalized by the partial slip, or Navier BCs models, which introduce the concept of slip length [4], being the distance across the boundary at which the linearly extrapolated fluid velocity would reach zero. For macroscopic objects, even when the slip length is nonzero, it is typically much smaller than the objects’ characteristic dimensions. As a result, the no-slip BCs provide a very good approximation in most cases. However, for micro- and nanoparticles, even small slip lengths can significantly influence particle dynamics. Consequently, partial slip BCs become important when dealing with systems approaching the molecular scale [5]. Partial slip BCs can influence the suspension’s viscosity [6, 7, 8, 9] and overall rheological behavior, especially in systems where interparticle interactions dominate [10]. Moreover, it has been suggested that the slip length may vary with the local shear rate, indicating a potential coupling between partial slip BCs and flow conditions [10, 11, 12].
Several experimental techniques have been developed to quantify slip lengths across various systems. One method involves tracking rough particles, where surface roughness mimics the effects of slip length [13, 14, 15]. Common experimental approaches for measuring slip length include surface force apparatus (SFA), atomic force microscopy (AFM), micro-particle image velocimetry (µ-PIV) [5], suspended microchannel resonator (SMR), dynamic quartz crystal microbalance (QCM-D), and hybrid graphene/silica nanochannel techniques [5]. Other indirect techniques such as pressure drop versus flow rate, drainage force measurements, and streaming potential analysis have also been employed [16]. Moreover, non-uniform slip conditions-such as those found in Janus particles with asymmetric surface properties—have been studied both experimentally [17] and theoretically [18, 19, 20, 21, 22].
In addition to experimental and theoretical studies, numerical methods have been used to investigate systems with partial slip BCs. These methods can be broadly categorized into explicit and implicit methods. Explicit methods directly resolve the fluid flow and include molecular dynamics (MD) [16, 23, 24], smoothed particle hydrodynamics (SPH) [25], smoothed dissipative particle dynamics (SDPD) [26], and dissipative particle dynamics (DPD) [27], among others. While effective, these methods are computationally demanding due to their need to discretize the fluid explicitly. In contrast, implicit methods do not solve for the flow field directly. Instead, they use the Green’s functions of the Stokes equations to compute the velocities of solute particles. These methods are limited to the study of suspensions in the Stokes regime. However, with this methods only the particles surfaces are discretized which reduces memory requirements considerably. Some implicit analytical methodologies have been also developed to include partial slip BCs [28, 29, 30, 9, 22].
This work presents a robust implicit method for modeling particles with partial slip BCs, capable of simulating complex and long-range interactions in particle suspensions. Additionally, the particles can feature non-homogeneous slip lengths along their surfaces. Our method employs a regularized boundary integral formulation to implement partial slip BCs and compute hydrodynamic tractions and particle velocities [31]. The approach achieves high accuracy, with only a few percent relative errors, even with a modest discretization per particle. Thus, it is appropriate for simulating large-scale particulate systems.
We validated our method by modeling spherical particles with uniform partial slip and no-slip conditions. We then study Janus particles, which exhibit asymmetric slip distributions over their surfaces, leading to complex anisotropic hydrodynamic behavior. Physically, a spherical particle with a uniform slip length tends to move along the direction of an applied force. However, in the case of non-uniform slip distributions, such as with Janus particles, the particle may move with an angle respect to the apply force and experience also a torque[32]. This effect naturally arises since the increase in the slip length reduces locally the fluid’s resistance, producing unbalance and asymmetric forces that influence the particle’s translational and rotational motion, even for spherical geometries [18].
The structure of the paper is the following: We begin in Section 2 by introducing the continuum formulation for modeling particles with partial slip BCs [33], including the definition of single-layer and double-layer operators. Section 3 details the discretization of the governing equations and the use of a preconditioned iterative solver to efficiently solve the Stokes problem. In Section 4, we validate the method by computing hydrodynamic drag on particles with both homogeneous and heterogeneous slip distributions over a wide range of slip lengths, comparing our numerical results with analytical predictions [6, 2, 34, 12]. Finally, we explore how slip length affects the relative viscosity of suspensions at different volume fractions, observing a decrease in viscosity with increasing slip length for a given volume fraction.
2 Continuum formulation
In this section we describe the formulation to model suspended rigid particles with partial slip BCs. We consider a collection of $M$ arbitrarily shaped particles, $\{\mathcal{B}_{m}\}_{m=1}^{M}$, immersed in a viscous dominated flow. At negligible Reynolds number the flow is governed by the Stokes equations
$-\boldsymbol{\nabla}p+\eta\boldsymbol{\nabla}^{2}\boldsymbol{v}$ $=\boldsymbol{0},$ (1)
$\boldsymbol{\nabla}\cdot\boldsymbol{v}$ $=0,$ (2)
where $\boldsymbol{v}$ and $p$ are the flow velocity and pressure and $\eta$ the fluid viscosity. The BCs on the particles surfaces are model as a slip condition
$\boldsymbol{v}(\boldsymbol{r})=\boldsymbol{u}_{m}+\boldsymbol{\omega}_{m}\times(\boldsymbol{r}-\boldsymbol{q}_{m})+\boldsymbol{u}_{s}(\boldsymbol{r})\text{ for $\boldsymbol{r}$ on }\partial\mathcal{B}_{m},$ (3)
where $\boldsymbol{u}_{m}$ and $\boldsymbol{\omega}_{m}$ are the linear and angular velocities of the particle $m$ around its tracking point (e.g. its center) $\boldsymbol{q}_{m}$ and $\boldsymbol{u}_{s}$ is the slip velocity. For $\boldsymbol{u}_{s}=\boldsymbol{0}$ the standard no-slip conditions are recovered. Instead, we use $\boldsymbol{u}_{s}$ to impose mixed Dirichlet-Neumann BCs on the flow velocity. We model the particle as impermeable particles, thus, the flow obeys the no-penetration boundary condition or, equivalently, the normal component of the slip vanishes, i.e. $\boldsymbol{u}_{s}(\boldsymbol{r})\cdot\boldsymbol{n}(\boldsymbol{r})=0$ with $\boldsymbol{n}(\boldsymbol{r})$ being the unit normal to the particle surface at $\boldsymbol{r}$. The tangential component of the slip is modeled with Navier BCs, i.e. the magnitude of the slip is proportional to the local hydrodynamic traction
$\xi(\boldsymbol{I}-\boldsymbol{n}\boldsymbol{n}^{T})\boldsymbol{u}_{s}=-(\boldsymbol{I}-\boldsymbol{n}\boldsymbol{n}^{T})\boldsymbol{\lambda},$ (4)
where $\boldsymbol{\lambda}$ is the traction exerted on the fluid and $\xi=\xi(\boldsymbol{r})$ the sliding friction coefficient that can vary over the particle surface [2, 3, 18]. We can write $\xi=\eta/\ell$, where $\eta$ is the fluid viscosity and $\ell=\ell(\boldsymbol{r})$ a characteristic length, the so-called slip length. In the limit $\xi\rightarrow\infty$ (equivalently $\ell=0$) the no-slip condition is recovered; the free-slip condition is achieved for $\xi=0$ ($\ell\rightarrow\infty$) [3]. Both conditions on the slip velocity can be combined into one equation
$\xi\boldsymbol{u}_{s}=-\boldsymbol{\mathcal{P}}\boldsymbol{\lambda}=-\left(\boldsymbol{I}-\boldsymbol{n}\boldsymbol{n}^{T}\right)\boldsymbol{\lambda},$ (5)
where $\boldsymbol{\mathcal{P}}=(\boldsymbol{I}-\boldsymbol{n}\boldsymbol{n}^{T})$ is the projector operator into the particle surface tangential space.
The equations are closed by the balance of force and torque, that is, the hydrodynamic traction on the particles balances the external forces, $\boldsymbol{f}_{m}$, and torques, $\boldsymbol{\tau}_{m}$, applied to them
$\int_{\partial\mathcal{B}_{m}}\boldsymbol{\lambda}(\boldsymbol{r})\,\mathrm{d}S_{\boldsymbol{r}}=\boldsymbol{f}_{m},$ (6)
$\int_{\partial\mathcal{B}_{m}}(\boldsymbol{r}-\boldsymbol{q}_{m})\times\boldsymbol{\lambda}(\boldsymbol{r})\,\mathrm{d}S_{\boldsymbol{r}}=\boldsymbol{\tau}_{m}.$ (7)
We make use of an integral formulation to solve the Stokes equations [35]. First, we introduce the the single layer operator, $\boldsymbol{\mathcal{M}}$, acting on the hydrodynamic traction $\boldsymbol{\lambda}$, and the Stokes double layer operator, $\boldsymbol{\mathcal{D}}$, acting on the fluid velocity
[…]
$\left(\boldsymbol{\mathcal{D}}\boldsymbol{v}\right)_{i}(\boldsymbol{x})$ $=\int_{\partial\mathcal{B}}\mathbb{D}_{ijk}(\boldsymbol{r}^{\prime},\boldsymbol{r}^{\prime\prime})n_{k}(\boldsymbol{y})v_{j}(\boldsymbol{y})\;\mathrm{d}S_{y},=\int_{\partial\mathcal{B}}-\frac{3}{4\pi}\frac{r_{i}r_{j}r_{k}}{r^{5}}n_{k}(\boldsymbol{y})v_{j}(\boldsymbol{y})\;\mathrm{d}S_{y},$ (9)
in an unbounded domain. Using these operators the double layer formulation for the fluid velocity in the particle surfaces is [35]
[…]
where $\boldsymbol{\mathcal{I}}$ is the identity operator. Replacing the slip condition, Eq. 3, one obtains [23, 33]
[…]
The equations (5)-(7) and (11) form a linear system for the unknown particle velocities, $\boldsymbol{u}_{p}$ and $\boldsymbol{\omega}_{p}$, hydrodynamic traction, $\boldsymbol{\lambda}$, and surface slip $\boldsymbol{u}_{s}$. In the next section we introduce a specific discretization and a method to solve the corresponding linear system.
3 Discretization
We discretize the surface of the particles with markers or blobs of radius $a$ with position $\boldsymbol{r}_{i}$. At each blob the surface normal, $\boldsymbol{n}_{i}=\boldsymbol{n}(\boldsymbol{r}_{i})$, is also defined, see Fig. 1. After the discretization, the balance of force and torque, Eqs. (6)-(7), become sums over the blobs of each body $m$
$\sum_{j\in\mathcal{B}_{m}}\boldsymbol{\lambda}_{j}$ $=\boldsymbol{f}_{m},$ (12)
$\sum_{j\in\mathcal{B}_{m}}(\boldsymbol{r}_{j}-\boldsymbol{q}_{m})\times\boldsymbol{\lambda}_{j}$ $=\boldsymbol{\tau}_{m}.$ (13)
We have included the quadrature weights into the traction, $\boldsymbol{\lambda}_{i}$, so they represent finite forces instead of density forces as in Sec. 2. The BCs for the partial slip, Eq. (5), are evaluated at each blob, that is
$\xi_{i}\boldsymbol{u}_{s,i}=-\boldsymbol{P}_{i}\boldsymbol{\lambda}_{i}=-w_{i}^{-1}\left(\boldsymbol{I}-\boldsymbol{n}_{i}\boldsymbol{n}_{i}^{T}\right)\boldsymbol{\lambda}_{i},$ (14)
where $w_{i}$, the quadrature weight of blob $i$, accounts for the fact that the Navier boundary condition is written in terms of density forces.
Figure 1: Examples of discretized spherical particles with different slip distributions: homogeneous no-slip, homogeneous full-slip and Janus slip distribution which we model with slip lengths $\ell/R=10^{-3}$ and $\ell/R=10^{3}$ respectively, where $R$ is the particle radius. The surface normals are denoted as n.
The partial slip condition, Eq. (11), is also evaluated on each blob $i$
$\sum_{j=1}^{N_{b}}\boldsymbol{M}_{ij}\boldsymbol{\lambda}_{j}$ $=\sum_{j=1}^{N_{b}}\left(\frac{1}{2}\boldsymbol{I}_{ij}+\boldsymbol{D}_{ij}\right)\left(\boldsymbol{u}_{m}+\boldsymbol{\omega}_{m}\times(\boldsymbol{r}_{j}-\boldsymbol{q}_{m})+\boldsymbol{u}_{s,j}\right),$ (15)
where $\boldsymbol{M}$ and $\boldsymbol{D}$ are discretizations of the single layer and double layer Stokes kernels and $\boldsymbol{I}_{ij}$ is a $3\times 3$ identity matrix for $i=j$ and zero otherwise.
The method uses a regularized version of the Stokes Green’s function, the Rotne-Prager mobility, which can be defined for non-overlapping blobs as [36, 37]
$\boldsymbol{M}_{ij}=\boldsymbol{M}(\boldsymbol{r}_{i},\boldsymbol{r}_{j})$ $=\left.\left[\boldsymbol{I}+\frac{a^{2}}{6}\boldsymbol{\nabla}^{2}_{\boldsymbol{r}^{\prime}}\right]\left[\boldsymbol{I}+\frac{a^{2}}{6}\boldsymbol{\nabla}^{2}_{\boldsymbol{r}^{\prime\prime}}\right]\mathbb{M}(\boldsymbol{r}^{\prime},\boldsymbol{r}^{\prime\prime})\right|_{\boldsymbol{r}^{\prime}=\boldsymbol{r}_{i}}^{\boldsymbol{r}^{\prime\prime}=\boldsymbol{r}_{j}}.$ (16)
For overlapping blobs it is possible to remove the singularity associated with the divergence of the Stokes kernels, see Ref. [37]. For reliable hydrodynamics interactions, is recommend a minimum distance of one blob radius $r_{o}=a$, such that the blobs do not overlap. For the double layer we use the same regularization procedure shown in Eq. (16).
Combining the four equations (12)-(15) gives the linear system
$\begin{bmatrix}\boldsymbol{M}&-\frac{1}{2}\boldsymbol{K}-\boldsymbol{D}\boldsymbol{K}&-\frac{1}{2}\boldsymbol{I}-\boldsymbol{D}\\ -\boldsymbol{K}^{T}&\boldsymbol{0}&\boldsymbol{0}\\ \boldsymbol{\xi}^{-1}\boldsymbol{P}&\boldsymbol{0}&\boldsymbol{I}\\ \end{bmatrix}\begin{bmatrix}\boldsymbol{\lambda}\\ \boldsymbol{U}\\ \boldsymbol{u}_{s}\\ \end{bmatrix}=\begin{bmatrix}\boldsymbol{0}\\ -\boldsymbol{F}\\ \boldsymbol{0}\\ \end{bmatrix},$ (17)
where the velocities and force-torque on the particles are collected on the vectors $\boldsymbol{U}=\{\boldsymbol{u}_{m},\boldsymbol{\omega}_{m}\}_{m=1}^{M}$ and $\boldsymbol{F}=\{\boldsymbol{f}_{m},\boldsymbol{\tau}_{m}\}_{m=1}^{M}$ and the unscripted vectors $\boldsymbol{\lambda}$ and $\boldsymbol{u}_{s}$ contain the traction and slip. The unscripted matrices are defined accordingly.
3.1 Numerical solver
Figure 2: Convergence of preconditioned GMRES to solve the mobility problem for a single particle discretized with a different number of blobs. The figure shows results for no-slip (left) and full-slip (right) BCs.
Figure 3: GMRES convergence for $\boldsymbol{M}$ particles, from 1 to 4096, forming a simple cubic lattice with a distance between first neighbors equal to two particle diameters. The particles are discretized with $N_{b}$ = 42 blobs.
The linear system (17) can be large for many-particle suspensions, therefore, using a direct solver is computationally unfeasible. Instead, we solve it with an iterative method. Since (17) is non-symmetric we employ a preconditioned GMRES solver which has been shown to be a very efficient for similar Stokes problems [38, 39].
To converge within a moderate number of iterations it is necessary to employ a preconditioner which approximates the inverse of (17). Since the preconditioner speeds up the convergence but it does not modify the solution of the linear system we can make a number of approximation so it can be applied efficiently. First, we note that in the continuum limit the action of the double layer on rigid body velocities is equivalent to applying half the identity operator [35]. Thus, on the preconditioner we approximate $\left(\frac{1}{2}\boldsymbol{I}+\boldsymbol{D}\right)\boldsymbol{K}\approx\boldsymbol{K}$. Second, in the mobility matrix we neglect the interaction between blobs belonging to different bodies, i.e. we replace the mobility matrix $\boldsymbol{M}$ by its block-diagonal approximation $\widetilde{\boldsymbol{M}}^{(pq)}=\delta_{pq}\boldsymbol{M}^{(pp)}$ where the superscripts $p$ and $q$ indicate the block of blobs belonging to body $p$ or $q$ respectively. Finally, we simply neglect the action of the double layer on the slip velocity. Applying the preconditioner derived with these approximations is equivalent to solve the system
$\begin{bmatrix}\widetilde{\boldsymbol{M}}&-\boldsymbol{K}&-\boldsymbol{I}\\ -\boldsymbol{K}^{T}&\boldsymbol{0}&\boldsymbol{0}\\ \xi^{-1}\boldsymbol{P}&\boldsymbol{0}&\boldsymbol{I}\\ \end{bmatrix}\begin{bmatrix}\boldsymbol{\lambda}\\ \boldsymbol{U}\\ \boldsymbol{u}_{s}\\ \end{bmatrix}=\begin{bmatrix}\boldsymbol{u}_{0}\\ -\boldsymbol{F}\\ \boldsymbol{B}\\ \end{bmatrix}.$ (18)
Note that it is necessary to include $\boldsymbol{u}_{0}$ and $\boldsymbol{B}$ in the right hand side because during the GMRES iterations those terms are not guarantee to be zero even if they are zero in the original Eq. (17). Since the preconditioner does not include many body interactions it can be solved easily for each body. The solution of the preconditioner is
$\boldsymbol{U}$ $=\left[\boldsymbol{K}^{T}\left(\widetilde{\boldsymbol{M}}+\xi^{-1}\boldsymbol{P}\right)^{-1}\boldsymbol{K}\right]^{-1}\left(\boldsymbol{F}-\boldsymbol{K}^{T}\left(\widetilde{\boldsymbol{M}}+\xi^{-1}\boldsymbol{P}\right)^{-1}(\boldsymbol{u}_{0}+\boldsymbol{B})\right),$ (19)
$\boldsymbol{\lambda}$ $=\left(\widetilde{\boldsymbol{M}}+\xi^{-1}\boldsymbol{P}\right)^{-1}\left[\boldsymbol{K}\boldsymbol{U}+\boldsymbol{u}_{0}+\boldsymbol{B}\right],$ (20)
$\boldsymbol{u}_{s}$ $=-\boldsymbol{\xi}^{-1}\boldsymbol{P}\boldsymbol{\lambda}+\boldsymbol{B}.$ (21)
We can see that the only expensive calculation to compute the action of the preconditioner correspond to computing the inverse $\left(\widetilde{\boldsymbol{M}}^{(pp)}+\xi^{-1}\boldsymbol{P}^{(pp)}\right)^{-1}$ for each body $p$. Since these matrices are defined for a single body they are not too big and their inverses can be computed with dense algebra methods, in particular, we use a Cholesky factorization.
We show in Fig. 2 the convergence of GMRES to solve the mobility problem for a single particle discretized with different number of blobs. We see that regardless of the discretization and the slip length the solver converges quickly, within a small number of iterations. In Fig. 3 we show the convergence for particles, discretized with $42$ blobs, forming a simple cubic lattice with distances betweeen first neighbors of two particle diameters. The figure shows that the convergence depends very weakly on the number of particles even for very large systems.
4 Numerical validation
In this section, we validate the scheme by computing the drag on spherical particles with different slip lengths and slip distributions. Each sphere, of geometric radius $R_{g}=1$, is discretized with $N_{b}$ blobs of radius $a$. We set the blob radius to $a=d/2$ where $d$ is the smallest distances between blobs so nearest neighbors touch each other. We will see that for a finite number of blobs the particle responses to the flow like a sphere with an effective hydrodynamic radius $R_{h}>R_{g}$. However $R_{h}$ converges to $R_{g}$ as the number of blobs increases. Figure 1 illustrates some of the discretized particles used in this section.
4.1 Mobility
First, we measure the effective hydrodynamic radius of a spherical particle with different resolutions to determine the scheme accuracy. Here we consider two cases, no-slip and full-slip BCs, which we implement using the slip lengths $\ell/R_{g}=10^{-6}$ and $\ell/R_{g}=10^{3}$ respectively. We apply a force or torque to the particles and define the effective radii as the translational hydrodynamic radius $R_{h}=\frac{f_{p}}{\alpha\pi\eta u_{p}}$, and the rotational radius as $R_{\tau}=\left(\frac{\tau_{p}}{8\pi\eta\omega_{p}}\right)^{1/3}$, where $\alpha=6$ or $\alpha=4$ for the no-slip and full-slip cases respectively. The measured radii are shown for several discretizations in Table 1. We note that for the full-slip BCs the rotational hydrodynamic radius is not well defined because no drag acts on a rotating particle. Therefore, in the full-slip case we only report the translational hydrodynamic radius. In all cases the values converge to $R_{g}$ with higher resolutions.
Number of blobs
$N_{b}$
Blob radius
$a/R_{g}$
No-Slip
$R_{h}/R_{g}$
No-Slip
$R_{\tau}/R_{g}$
Full-Slip
$R_{h}/R_{g}$
12 0.5257 1.0819 1.0826 1.0345
42 0.2733 1.0321 1.0338 1.0046
162 0.1380 1.0086 1.0152 0.9942
642 0.06914 1.0020 1.0072 0.9944
2562 0.03459 1.0003 1.0035 0.9964
Table 1: Effective translational and rotational radii for spheres discretized with different number of blobs. Results for sphere with homogeneous no-slip and full-slip BCs.
Table 1 shows that both the translational and rotational hydrodynamic radius are close for all discretizations. In practical applications, if we wanted to study particles of radius $R$ we would set the hydrodynamic radius to the target radius of the particles under study $R_{h}=R$ (thus, $R_{g}<R$ for finite resolutions). This allows us to get results accurate up to a few percent using coarse discretizations. For complex shaped particles one can compute the particle mobility to high accuracy and then optimize a low resolution grid to match the high accuracy computed mobility [40, 39]
4.2 Drag
As second test to validate our partial slip BCs, we solve a mobility problem to compute the drag force $f_{d}$ on a sphere with homogeneous slip but different slip length values $\ell$. We show in Fig. 4 the numerical results for two discretizations with $N_{b}=42$ and $642$ respectively and the analytical drag force $f_{d}=6\pi\eta R\left(\frac{1+2\ell/R}{1+3\ell/R}\right)u$ [2, 41]. We normalize our numerical results with the hydrodynamic radius, $R_{h}$, computed in the no-slip case. For the coarser resolution with 42 blobs, the numerical error is below $3\%$ for all slip lengths while for the fine resolution with 642 blobs is below $1\%$. This showcases that low resolutions can be used to get accurate results up to a few percent error.
Figure 4: Drag force $f_{d}$ on a sphere with different slip lengths, $\ell$, ranging from no-slip to full-slip. The curve represents the analytical result and the symbols the numerical results for two discretizations with $N_{b}=42$ and 642 blobs respectively. The numerical results are normalized with the hydrodynamic radius, $R_{h}$, computed in the no-slip case ($\ell/R\approx 10^{-4}$).
4.3 Janus particles
In the previous validations we tested spherical particles across a full range of slip conditions, from no-slip to full-slip, but always with an homogeneous slip. However, the proposed scheme allows for definition of heterogeneous slip conditions over the surface of the particle, such can occur in decorated patchy nanoparticles or Janus particles. Here, we measure the drag force and torque on a Janus particle following the work of Sun et al. [29]. On the Janus particle we apply non-homogeneous BCs: one half is set to full-slip ($\ell/R_{g}=10^{4}$) while the other half is set to no-slip ($\ell/R_{g}=10^{-4}$) and an intermediate region is assigned a slip length that is the average of these two values, see Fig. 5. Instead of solving the mobility problem, i.e. applying forces and torques to compute the body velocities, we solve a resistance problem. We set the particle velocities to $\boldsymbol{u}=(0,0,1)$ and $\boldsymbol{\omega}=\boldsymbol{0}$ and compute the resulting force-torque $\boldsymbol{F}$ for different orientations respect to the direction of motion as shown in Fig. 5 [29].
Figure 5: Janus particles have one half (colored yellow) representing a no-slip surface, while the other half (colored purple) represents a full-slip surface. The interphase line region is an average slip length. The scheme illustrates examples of Janus particles aligned along the P. axis at three angles: 0, 45, and 90 degrees from the horizontal x-axis. In all cases, the applied force acts along the horizontal x-axis.
Figure 6 presents the drag force and torque for a Janus particle discretized with different number of blobs. The results for $f_{z}$ converge to the analytical result of Sadhal and Johnson for $\theta=0\text{ and }\pi$, see 6b [42]. For other angles or force-torque components our results show self-convergence. We should mention that our results also agree to within $1\%$ to the results’ of Sun et al. [29] (not shown) which have an error of $1\%$ respect the analytical result of Sadhal and Johnson.
Figure 6: Force ($f_{z},f_{x}$) and torque ($\tau_{y}$) components for spherical Janus particle as functions of the angle $\theta$. Results for discretization with different numbers of blobs, $N_{b}$. The analytical result of Sahdal for $f_{z}$ at $\theta=0,\pi$ is shown as a red point in panel (b) [42].
4.4 Suspension viscosity
In this section we measure the high frequency viscosity in a suspension of spherical particles with homogeneous slip length and we explore how the suspension viscosity changes with volume fraction and slip length. We study suspensions with 250 particles in a triply periodic domain of size $L$ which we vary to set the desired volume fraction $\phi$. In this section we use particles discretized with $N_{b}=162$ blobs. To measure the high frequency shear viscosity we follow the following protocol. First, we generate 100 random configurations using a Monte Carlo code. Then, for each configuration we apply a sinusoidal background flow
$\boldsymbol{v}_{0}(x,y,z)=v_{0}\sin\left(\frac{2\pi z}{L}\right)\boldsymbol{e}_{x},$ (22)
and we measure the magnitude of the flow velocity profile $v_{x}$. Then, we extract the viscosity from the flow magnitude as
$\eta_{\text{susp}}=\frac{v_{x}}{v_{0}}.$ (23)
However, to obtain accurate results some care it is necessary in applying the background flow and in measuring the resulting flow. As the Stokes equations are linear the background flow can be included in the right hand side of the slip equation, that is
$\boldsymbol{M}\boldsymbol{\lambda}-\left(\frac{1}{2}\boldsymbol{I}+\boldsymbol{D}\right)\boldsymbol{K}\boldsymbol{U}-\left(\frac{1}{2}\boldsymbol{I}+\boldsymbol{D}\right)\boldsymbol{u}_{s}=-\boldsymbol{v}_{0}.$ (24)
However, we obtained results with large numerical errors using this straightforward approach.
[…]
Our strategy to reduce the numerical error is to add a correction term to the background flow that vanishes in the continuum limit (number of blobs $N_{b}\rightarrow\infty$) but that guarantees that particles are advected with the correct velocity for constant background flows for all resolutions. These two conditions are satisfied by adding and subtracting a rigid body contribution to the background flow, so in the right hand side of Eq. (24) we use the effective flow
$\boldsymbol{v}_{\text{RHS}}=\boldsymbol{v}_{0}-\boldsymbol{K}\left(\frac{1}{N_{b}}\boldsymbol{K}^{T}\boldsymbol{v}_{0}\right)+\left(\frac{1}{2}\boldsymbol{I}+\boldsymbol{D}\right)\boldsymbol{K}\left(\frac{1}{N_{b}}\boldsymbol{K}^{T}\boldsymbol{v}_{0}\right).$ (25)
In the continuum limit, the last two terms in (25) cancel out, thus the continuum problem is not modified. In the discrete case, for constant background flow, the first two terms cancel exactly and the last term balances the term $\left(\frac{1}{2}\boldsymbol{I}+\boldsymbol{D}\right)\boldsymbol{K}\boldsymbol{U}$ in (24) which guarantees that particles are advected with the background velocity to machine precision. Using this background flow it is possible to obtain accurate viscosity results as we will show in a moment.
The second detail to extract the viscosity with higher accuracy is to measure the flow velocity correctly. Since we solve the Stokes equations with an implicit method, we do not have direct access to the flow (although it can be computed). Therefore, we measure the average particle velocities profile, $\langle u_{x}\rangle(z)$, and use the Faxén law to extract the ambient velocity magnitude from
$\langle u_{x}\rangle=\left(\boldsymbol{I}+\frac{R_{h}^{2}}{6}\boldsymbol{\nabla}^{2}\right)v_{x}\cos\left(\frac{2\pi z}{L}\right).$ (26)
The flow magnitude, $v_{x}$, extracted from this equation is used in Eq. (23) to compute the suspension velocity.
Figure 7: High frequency viscosity for suspension of spherical particles. (a) Relative viscosity increase at low volume fraction ($\phi=0.01$) versus slip length. (b) Relative viscosity versus volume fraction for several slip lengths. Symbols represent numerical results, continuous lines the theoretical prediction of Ref. [12] and dashed lines no-slip Stokesian results [34].
In the left panel of Fig. 7 we show the viscosity at low volume fraction, $\phi=0.01$, versus the slip length. We compare the numerical results with analytical result from Happel and Brenner [2]
$\frac{\eta_{\text{eff}}}{\eta}=1+\frac{5}{2}\cdot\frac{1+2\ell/R}{1+5\ell/R}\phi,$ (27)
which go from the no-slip limit (Einstein’s limit) to the full-slip limit [6, 2]. We observe a good agreement between numerical and theoretical results for all the slip lengths considered.
Finally, in the right panel of Fig. 7 we show the suspension viscosity versus volume fraction for several slip lengths. The no-slip results agree well with the Stokesian dynamics results of Sierou and Brady [34], while the viscosity grows slower for larger slip lengths as expected. The viscosity decreases with the slip length until $\ell/R\approx 10$. Further increases in the slip length do not reduce the viscosity any lower. We also compare our numerical results with an approximated theory derived from a Stokes–Darcy model with partial slip [12]. For small slip lengths (e.g., $\ell\leq 0.1$), the numerical results show good agreement with the analytical predictions. However, for larger slip lengths, the theory proposed in Ref. [12] consistently overpredicts the effective viscosity. This overprediction occurs across all volume fractions, as the theory’s prediction for low concentrations of free-slip particles, $\eta_{\text{eff}}/\eta=1+1.6\phi$, exceeds the exact prediction, to linear order, given by Eq. (27). Nevertheless, the theory successfully captures the overall trend observed in the numerical results.
5 Conclusions
Deviations from the no-slip BCs can be modeled using partial slip BCs [3, 4]. This paper introduces a numerical method for simulating particle suspensions with partial slip BCs. Our approach is based on a regularized boundary integral method. It is easy to implement, accurate within a few percent, and capable of simulating large-scale suspensions. The scheme employs a regularized double-layer formulation to implement the partial slip BCs [23, 33]. Only the particle surfaces are discretized within this framework, significantly reducing the algorithm’s memory requirements. To solve the Stokes equations, we use a preconditioned GMRES solver. With a simple block-diagonal preconditioner, the number of iterations required to solve the Stokes equations depends weakly on the number of particles or the resolution used (see Figs. 23).
We have validated the accuracy of our scheme across a range of slip lengths, from full-slip to no-slip BCs. In all cases, our results remained accurate within a few percent, even with resolutions as low as 42 blobs per particle. Additionally, we have reproduced the drag forces and torques on Janus particles. We have also measured the high-frequency viscosity in suspensions of particles with homogeneous slip distributions. At low volume fractions, our viscosity results are in good agreement with analytical predictions [2]. Over a broad range of volume fractions, our results are consistent with Stokesian dynamics simulations in the limit of small slip lengths ($\ell/R\approx 0.01$) while, as the slip length increases, the viscosity decreases as expected. We found that the viscosity decreases with the slip length up to $\ell/R=10$, beyond which it becomes independent of the slip length.
The scheme presented in this work is flexible enough to model inhomogeneous slip length distributions, incorporating for instance, regular and irregular patterns over the surfaces, that have potential applications in nanotechnology and biomedicine [43, 44, 45]. The extension of this work to Brownian suspensions would be an important contribution. However, how to generate Brownian velocities when the double layer operator, which is non-symmetric and thus not positive definite, appears in Eq. (11) remains an open problem. Our implementation of the method is publicly available as free software at https://github.com/stochasticHydroTools/RigidMultiblobsWall.
Acknowledgments
We acknowledge funding by the Basque Government through the BERC 2022-2025 program and by the Ministry of Science, Innovation and Universities: BCAM Severo Ochoa accreditation CEX2021-001142-S/MICIN/AEI/10.13039/501100011033. The Spanish State Research Agency through the project PID2020-117080RB-C55 funded by (AEI/FEDER, UE) with acronym COMPU-NANO-HYDRO.
Declaration of generative AI and AI-assisted technologies in the writing process:
During the preparation of this work the author(s) used Chat GPT and Grammarly in order to check the grammar and readability of the text. After using this tool/service, the author(s) reviewed and edited the content as needed and take(s) full responsibility for the content of the publication.